# CasTex: Cascaded Text-to-Texture Synthesis via Explicit Texture Maps and Physically-Based Shading

### v1

@@ 1 Introduction
The creation of 3D assets is a crucial part of modern computer graphics applications, with demand being supported by such industries as movie production, industrial design, and computer gaming, to name a few. High-quality assets should simultaneously meet aesthetic requirements and technical production requirements (i.e., high-quality mesh topology and limited polygon count). As a result, designing a 3D model is a labor-intensive creative task that involves multiple skills and is hard to automate due to the lack of high-quality data and occasionally difficulty in formalizing target criteria.
@@
At the same time, the rapid development of generative models in the visual domain has already tremendously impacted various creative domains and even extends to 3D graphics. The successes in 3D reconstruction [25, 1, 17] and foundation text-to-image model [33, 32] led fully-automated end-to-end 3d asset creation prototypes [29, 23, 20]. Such solutions use the prior knowledge acquired by the diffusion model to generate 3D assets meeting aesthetic requirements. However, the resulting assets come in a volumetric representation and are often incompatible with modern graphics engines. Even with conversion tools for extracting meshes from volumetric representation, the assets end up having overly detailed geometry with occasional high-level geometric inconsistencies (e.g., multiple faces, as in the Janus effect). The aforementioned limitations may be critical when the technical requirements come into play.
Texture synthesis is a more modest attempt to automate 3D asset creation, aiming to create a realistic texture for a given 3D shape. The task assumes that all of the technical requirements regarding the model geometry are already addressed by an artist. To meet modern industry standards, in our work, we specifically consider the synthesis of physically based textures [4]. Such textures present a versatile instrument able to capture a diversity of materials and create realistically looking assets within varying lighting scenarios.
@@
Some early texture synthesis approaches involved the back-projection of synthetic images on 3D shapes. However, such approaches are incompatible with PBR textures as they only allow the capture of a fraction of material properties through RGB colors. As an alternative, we consider an optimization based approach that allows capturing the material properties with back-propagation through the rendering engine. In particular, we continue the line of work studying various applications of Score Distillation Sampling [29].
@@
We aimed to design a zero-shot texture-synthesis solution without relying on auxiliary 3D data and foundation model fine-tuning. However, we note that the texture synthesis setup unveils the limitations of Score Distillation Sampling applied in conjunction with modern diffusion models. Our observations go in line with the results of [41]. We continue their analysis and attribute the effect to the specifics of latent diffusion models.
@@
As an alternative, we consider cascaded diffusion models that guide the images in RGB space rather than latent space. Our solution follows the pipeline of cascaded models. First, we generate a coarse texture using SDS guidance with a low-resolution model. Then we refine the texture using SDS guidance with a super-resolution model. Below, we summarize our main contributions.
@@
  • We study the limitations of Score Distillation Sampling when applied to latent diffusion models. Our analysis indicates that the optimal solutions of the algorithm include images outside of the natural image domain.
@@
  • Motivated by our analysis, we introduce a texture synthesis approach using Score Distillation Sampling with cascaded diffusion models. The proposed approach mimicks the diffusion model architecture and acts in two stages. The first stage generates a coarse texture, while the second stage improves overall quality of the texture obtained at the first stage.
@@
  • We show that our approach outperforms state-of-the-art optimization based texture synthesis solutions on the Objaverse dataset. Score Distillation Sampling in pixel space allows omitting additional regularizers required in previous works. Additionally, we report the scaling performance of our method with the diffusion model size, the impact of lighting conditions used during training, and the role of physically based textures.
@@ 2 Related Work
Generating high-quality textures for 3D objects is a long-standing challenge in computer graphics and machine learning. Early approaches are based on category-specific generative models, such as generative adversarial networks (GANs) like Texture Fields [28], SPSG [11], LTG [42], and Texturify [36]. These methods often struggle with limited generalization and required extensive manual adjustments. Recent advances in large-scale text-to-image diffusion models [32, 33] transform texture synthesis, allowing more flexible and realistic results. However, different approaches vary in how they leverage diffusion models to incorporate 3D information and ensure global texture consistency.
@@
An alternative strategy to address inconsistency is to introduce additional structures to enforce coherence between views. SyncMVD [21] tackles this by generating each view independently while employing a synchronization module that aligns outputs in every denoising step of the diffusion model. Similarly, TexFusion [5] suggest a module that aggregates multi-view object predictions into a unified latent texture representation, iteratively refining it throughout the diffusion process to ensure consistency across the entire object.
@@
Another approach, inspired by MVDream [34], leverages consistent multi-view image generation as an initialization step for texture synthesis. Meta 3D TextureGen [2] follows this idea by generating four consistent views, which are projected onto a 3D surface to form an initial texture for subsequent refinement in UV space. MVPaint [8] extends this pipeline by introducing a spatial-aware 3D inpainting stage between projection and UV refinement, further enhancing coherence across different object views.
@@
Optimization-based methods. Optimization-based techniques refine a texture by optimizing an objective function that evaluates rendered views. Early approaches such as CLIP-Mesh [26] and Text2Mesh [24] use differentiable rendering and CLIP [30] as a guidance signal to iteratively update texture parameters. However, these methods often produce blurry textures due to the lack of high-frequency details.
@@
More recent works incorporate diffusion models to improve texture optimization. Latent-Paint [23], Fantasia3D [7], and Paint-it [41] combine differentiable rendering with the Score Distillation Sampling (SDS) technique [29], distilling knowledge from a diffusion model to refine textures. While this approach enhances realism, it often suffers from color oversaturation and high-frequency artifacts. FlashTex[13] mitigates some of these issues by employing a fine-tuned multi-view model to generate an initial textured view of the object, followed by an optimization stage using L2-loss and SDS for refinement. In this work, we continue the line of work employing SDS. However, we attribute some of the artifacts to the underlying latent diffusion model and propose using a pixel-space model. Aiming at physically-plausible textures, we use a text-to-image model pretrained on natural images and resort from using synthetic data for multi-view model adaptation.
@@
PBR texture generation. Most of the works discussed above generate only RGB, yielding limited photo-realism. The works of Fantasia3D, Paint-it, and FlashTex propose to synthesize more advanced physically based textures [4], which are able to mimic different materials and look realistic under different lighting. We asses a similar approach to generation of PBR textures with a different generative model.
3 Preliminaries
In our work, we rely on a pretrained text-to-image model to generate textures. To apply a model designed for image generation to a different task, we use Score Distillation Sampling (SDS). Next, we describe the algorithm and the rendering algorithm required to apply it.
3.1 Score Distillation Sampling for Gradient Based Generation
SDS gained popularity in text-to-3D setup. In [29], the authors consider a differentiable rendering mechanism that produces a random view of a 3D model. In particular, they parameterize the model with a volumetric radiance field and render it using random lighting and camera position. SDS leverages a pretrained text-to-image diffusion model to guide the model parameters with gradient descent. Crucially, the mild assumptions on the rendering algorithm allow one to apply the method to other differentiable image generation mechanisms. In later works, Score-Distillation Sampling was successfully applied to other rendering functions and image parameterizations such as Gaussian Splatting [40, 9, 37], vector graphics [39, 14], and image rasterization with physically-based shading [20, 7, 41].
@@ 3.1 Score Distillation Sampling for Gradient Based Generation
SDS defines gradients used to update the parameter $\theta\in\Theta$. Assume a mapping $x_{0}=g(\theta,z)$ that takes the parameter and a random variable $c$ and produces an image $x_{0}$. Given a diffusion process with noisy images $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ at time $t$ and scale-parameter $\bar{\alpha}_{t}$, the approach leverages a pretrained diffusion model $\epsilon_{\phi}(x_{t};y,t)$ to define a pseudo-objective used to update $x_{0}$. [^1] Throughout the paper, we assume that the diffusion model $\epsilon_{\phi}$ incorporates the condition $y$ through classifier-free guidance. The pseudo-objective $\hat{x}_{0}^{t}=\tfrac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\phi}(x_{t},t,y)}{\sqrt{\bar{\alpha}_{t}}}$ estimates the appearance of the rendered image $x_{0}$ based on a noisy sample $x_{t}$ with a single step diffusion model prediction.
The resulting gradient is estimated as follows:
@@
where $w(t)$ is an adjustable scalar weight parameter. Note that [29] formulate the method in noise prediction $\epsilon$-terms, while we introduce the equivalent formulation in $x_{0}$-terms. On the one hand, the gradient is essentially the mean-squared loss gradient $\mathbb{E}\left[w(t)\|x_{0}-\hat{x}_{0}^{t}\|^{2}\right]$ with the omitted dependence of $\hat{x}_{0}^{t}$ on $\theta$. On the other hand, the analysis in [29] demonstrates that the procedure minimizes $\mathcal{L}_{SDS}(\theta)=\mathbb{E}_{t,c}\operatorname{KL}\left(q(x_{t}\mid t)\|p_{\phi}(x_{t}\mid y,t)\right)$ using the score $\epsilon_{\phi}(x_{t};y,t)=\nabla_{x_{t}}\log p_{\phi}(x_{t}\mid y,t)$.
Originally, SDS was applied to a pixel-space diffusion model that generates low-resolution images. However, pretrained pixel-space models are less common, and, as a result, applications of the method were primarily focused on latent diffusion models (LDMs). These models approximate the distribution of the low-dimensional embeddings of a high-resolution images rather than the images itself. Crucially, the mapping $\text{enc}(\cdot)$ that produces the embedding is differentiable, as a result, can be considered as a component of a function generating an image:
@@
The potential advantage of latent diffusion models is the ability to guide high-resolution images compared to pixel-space models. At the same time, it requires additional resources to run forward and backward passes of the encoder network. In Section 4, we analyze other implications of applying SDS in latent space rather than pixel space.
@@ 3.2 Differentiable Physically Based Shading
Our texture synthesis approach applies SDS in a setting, where $g(\theta,z)$ renders an image of a fixed 3D model given a set of texture parameters $\theta$ and a random camera position and scene lighting specified by $z$. For rendering, we follow the pipeline proposed by [27].
In particular, we take a differentiable rasterization algorithm based on the Nvdiffrast library by [19]. For textures, we adopt the Disney-physically based material model from [4], allowing us to incorporate the obtained textures into popular graphics engines without further modifications. The texture $\theta=(\theta_{d},\theta_{r},\theta_{m},\theta_{n})$ consists of a diffuse component $\theta_{d}\in\mathbb{R}^{h\times w\times 3}$, a roughness map $\theta_{r}$ parameterizing a specular lobe and a metalness factor $\theta_{m}$ used to interpolate between plastic and metallic appearance. Finally, a normal map $\theta_{n}$ stores surface normal information, allowing to simulate additional geometric detail even within low-polygon models.
To model lighting we incorporate cube environment map with a split sum approximation for the outgoing radiance integral approximation [16].
Figure 1: Overall pipeline. Our method consists of two stages. On the first stage, given a 3D mesh and a prompt our method employs a differentiable rendering pipeline to generate random views of the model under various lighting. We update the randomly initialized model texture using Score Distillation Sampling (SDS). On the second stage, refine the texture from the first stage using SDS with a super-resolution diffusion model. For the same view and lighting, we render two frames: a frame with the fixed texture taken from the first step and a frame with a current texture. Using the former frame as the condition, we back-propagate the SDS gradients through the latter frame.
Figure 2: Optimization based image generation for a toy single-mode diffusion model. We study the effect of image parameterization and the difference between pixel and latent space diffusion models. Score Distillation Sampling (SDS) finds the distribution mode for both latent and pixel space diffusion models. Top row: the pixel space model achieves near-perfect reconstruction quality measured by PSNR. The latent code $\text{enc}(\theta_{\text{RGB}})$ is also close to the mode and leads to high quality reconstruction. However, for the latent model the optimized parameter $\theta_{\text{RGB}}$ fails to approximate the original image. Bottom row: The implicit regularization proposed in [41] significantly improves the results for SDS for the latent diffusion model, but still falls short of the image reconstruction given by the decoder or pixel space diffusion model.
4 Methodology
In the following section, we start with a motivating example and then describe the proposed method. The overall scheme is illustrated in Figure 1.
4.1 Limitations of Score Distillation Sampling for Latent Diffusion Models
In this section, for latent diffusion models, we argue that the procedure is prone to producing samples with artifacts due to the ill-posed nature of the task.

### v2

@@ 1 Introduction
The creation of 3D assets is a central component of modern computer graphics, with demand driven by industries such as film production, industrial design, and computer gaming. High-quality assets should simultaneously meet aesthetic requirements and technical production requirements (e.g., clean mesh topology and limited polygon counts). As a result, 3D modeling remains a labor-intensive creative task that requires multiple expert skills and is difficult to automate due to the scarcity of high-quality data and the challenge of formalizing target criteria.
@@
Meanwhile, the rapid progress of generative models in the visual domain has already transformed many creative domains and is now extending into 3D graphics. Advances in 3D reconstruction [23, 1, 15] and text-to-image foundation models [31, 30] have enabled early propotypes for fully automated, end-to-end 3d asset generation [27, 21, 18]. These systems leverage the prior knowledge acquired by the diffusion model to generate visually appealing 3D content. However, the resulting assets come in a volumetric representation and are often incompatible with modern graphics engines. Even when volumetric outputs are converted into meshes, the geometry tends to be excessively detailed and may contain high-level structural errors (e.g., multiple faces as in the Janus effect). The limitations mentioned above may be critical when technical requirements must be met.
Texture synthesis offers a more targeted and practical path toward automating parts of the 3D asset pipeline by generating realistic surface appearance for a given shape. In this work, we assume that the geometry is provided by an artist and focus on synthesizing physically based textures [4], which allow for expressive material modeling and realistic rendering under diverse lighting conditions.
@@
Early texture synthesis methods relied on back-projecting synthetic images onto 3D surfaces. However, such approaches are incompatible with PBR materials because RGB images capture only a limited subset of material properties through RGB colors. As an alternative, we consider an optimization-based framework that enables back-propagation through a differentiable renderer to recover full material parameters. In particular, we build on a line of recent work exploring Score Distillation Sampling (SDS) [27] for optimization-based generation.
@@
Our goal is to develop a zero-shot texture-synthesis solution that avoids reliance on auxiliary 3D data and foundation model fine-tuning. However, we observe that the texture synthesis setup unveils the limitations of Score Distillation Sampling applied with modern diffusion models. Consistent with [36], we find that SDS often drives optimization toward images that drift outside the natural image manifold. We expand upon this observation and attribute the effect to inherent properties of latent diffusion models.
@@
Motivated by this analysis, we explore cascaded diffusion models, which operate in RGB space rather than latent space. Our method follows the cascaded-model structure: we first generate a coarse texture using SDS guided by a low-resolution model, then refine it using SDS with a super-resolution model. In the following, we summarize our main contributions.
@@
  • We study the limitations of Score Distillation Sampling when applied to latent diffusion models. We show that the corresponding optimization procedure tends to converge to images outside the natural image domain.
@@
  • Based on these insights, we introduce CasTex, a texture-synthesis approach using Score Distillation Sampling with cascaded pixel-space diffusion models. The proposed method mimics the architecture of a diffusion model and acts in two stages. The first stage generates a coarse texture, while the second stage improves the overall quality of the texture obtained in the first stage.
@@
  • We demonstrate that CasTex outperforms state-of-the-art optimization-based texture synthesis solutions on the Objaverse dataset. Score Distillation Sampling in pixel space allows omitting additional regularizers required in previous works and enables efficient, high-resolution texture generation. Additionally, we report the scaling performance of our method with the diffusion model size, the impact of lighting conditions used during training, and the role of physically-based textures.
@@ 2 Related Work
Generating high-quality textures for 3D objects is a long-standing challenge in computer graphics and machine learning. Early approaches are based on category-specific generative models, such as generative adversarial networks (GANs) like Texture Fields [26], SPSG [10], LTG [37], and Texturify [34]. These methods often struggle with limited generalization and required extensive manual adjustments. Recent advances in large-scale text-to-image diffusion models [30, 31] have transformed texture synthesis, allowing more flexible and realistic results. However, different approaches vary in how they leverage diffusion models to incorporate 3D information and ensure texture consistency.
@@
An alternative strategy to address inconsistency is to introduce additional structures to enforce coherence between views. SyncMVD [20] tackles this by generating each view independently while employing a synchronization module that aligns outputs in every denoising step of the diffusion model. Similarly, TexFusion [5] suggest a module that aggregates multiview object predictions into a unified latent texture representation, iteratively refining it throughout the diffusion process to ensure consistency across the object.
@@
Another approach, inspired by MVDream [32], leverages consistent multiview image generation as an initialization step for texture synthesis. Meta 3D TextureGen [2] follows this idea by generating four consistent views, which are projected onto a 3D surface to form an initial texture for subsequent refinement in UV space. MVPaint [8] extends this pipeline by introducing a spatially-aware 3D inpainting stage between projection and UV refinement, further enhancing coherence across different object views.
@@
Optimization-based methods. Optimization-based techniques refine a texture by optimizing an objective function that evaluates rendered views. Early approaches such as CLIP-Mesh [24] and Text2Mesh [22] use differentiable rendering with CLIP [28] as a guidance signal to iteratively update texture parameters. While these methods are simple and training-free, they tend to produce blurred textures lacking high-frequency detail.
@@
More recent work incorporate diffusion models to improve texture optimization. Latent-Paint [21], Fantasia3D [7], and Paint-it [36] combine differentiable rendering with Score Distillation Sampling (SDS) technique [27], distilling knowledge from a diffusion model to refine textures. Although this approach enhances realism, it often suffers from color oversaturation and high-frequency artifacts. FlashTex[12] mitigates some of these issues by using a fine-tuned multiview model to generate an initial textured view of the object, followed by an optimization stage using L2-loss and SDS for refinement.
In this work, we continue the line of work employing SDS. However, we attribute some of the artifacts to the underlying latent diffusion model and propose using a pixel space model. Aiming at physically plausible textures, we use a text-to-image model pretrained on natural images and resort from using synthetic data for multiview model adaptation.
@@
PBR texture generation. Most of the works discussed above generate only RGB, limiting physical realism. Fantasia3D [7], Paint-it [36], and FlashTex [12] propose to synthesize more advanced physically based textures [4], which are able to mimic different materials and look realistic under different lighting. We pursue a similar goal but employ a different class of generative models and demonstrate the advantages of SDS in pixel space.
Figure 1: Overall pipeline. Our method consists of two stages. In the first stage, given a 3D mesh and a text prompt, we use a differentiable rendering pipeline to generate random views of the model under varying lighting conditions. The texture, initialized randomly, is optimized using the SDS gradient $\nabla_{\theta}\mathcal{L}_{\text{SDS}}(\theta)$ (Eq. 1). In the second stage, we refine the texture produced in stage one using SDS with a super-resolution diffusion model. For each camera view and lighting setup, we render two images: one using the fixed texture from the first stage, and one using the current texture being optimized. We denote the SDS gradient adapted to the super-resolution model as $\nabla_{\theta}\mathcal{L}_{SR-SDS}$. Using the former image as a conditioning input, we backpropagate the SR-SDS gradients through the latter image.
3 Method
We leverage a pretrained text-to-image diffusion model to synthesize high-quality textures. Because such models are trained for unconditional image generation rather than texture optimization, we employ Score Distillation Sampling (SDS) to obtain gradients compatible with our differentiable rendering pipeline. Specifically, we couple SDS with a cascaded diffusion model to refine textures across multiple resolutions. The overall architecture is illustrated in Figure 1.
3.1 Preliminaries on Score Distillation Sampling for Gradient Based Generation
@@
SDS defines gradients used to update the parameter $\theta\in\Theta$. Let $x_{0}=g(\theta,c)$ denote the mapping from the parameter $\theta$ and a random variable $c$ to a rendered image $x_{0}$. Given a diffusion process in which noise images are generated as $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ at timestep $t$ with noise scale $\bar{\alpha}_{t}$, the method leverages a pretrained diffusion model $\epsilon_{\phi}(x_{t};y,t)$ to define a pseudo-objective that guides updates to $x_{0}$. [^1] Throughout the paper, we assume that the diffusion model $\epsilon_{\phi}$ incorporates the conditioning $y$ via classifier-free guidance. The pseudo-objective $\hat{x}_{0}^{t}=\tfrac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\phi}(x_{t},t,y)}{\sqrt{\bar{\alpha}_{t}}}$ provides a single-step estimate of the clean image $x_{0}$ from the noisy sample $x_{t}$ using the diffusion model’s prediction.
The resulting gradient is estimated as
@@
where $w(t)$ is a scalar weight parameter. While [27] express SDS in terms of noise prediction $\epsilon$, we present an equivalent formulation in the $x_{0}$-domain. Viewed from one perspective, this gradient has the form of a mean-squared-error objective, $\mathbb{E}\left[w(t)\|x_{0}-\hat{x}_{0}^{t}\|^{2}\right]$, where the depencence of $\hat{x}_{0}^{t}$ on $\theta$ is ignored when computing the gradient. From another perspective, the analysis in [27] shows that SDS minimizes the divergence $\mathcal{L}_{SDS}(\theta)=\mathbb{E}_{t,c}\operatorname{KL}\left(q(x_{t}\mid t)\|p_{\phi}(x_{t}\mid y,t)\right)$ where the pretrained diffusion model provides the score $\epsilon_{\phi}(x_{t};y,t)=\nabla_{x_{t}}\log p_{\phi}(x_{t}\mid y,t)$.
Originally, SDS was applied to pixel-space diffusion models that generate low-resolution images. However, pretrained pixel-space models are relatively rare, and consequently most applications of SDS have focused on latent diffusion models (LDMs). These models approximate the distribution of low-dimensional latent embeddings of high-resolution images rather than the images themselves. Importantly, the encoder mapping $\text{enc}(\cdot)$ that produces these embeddings is differentiable and can therefore be incorporated directly into the image-generation pipeline:
@@
A key advantage of LDMs is their ability to guide the synthesis of high-resolution images more efficiently than pixel-space models. However, this comes at the cost of additional computation due to the encoder’s forward and backward passes. In Section 4, we further examine the implications of applying SDS in latent space rather than pixel space.
Figure 2: Score distillation sampling applied to a super-resolution diffusion model. The horizontal axis indicates optimization progress over time. Top row: the same image $g(\theta)$ is used both as a diffusion-model input $x$ and as the conditioning image $x_{cond}$ for super-resolution. This setup is unstable and causes the optimization to drift away from the original image. Bottom row: the same procedure, but with fixed super-resolution condition $x_{cond}=g(\theta_{0})$. The fixed condition anchors the optimization dynamics, preserving the structure of the initial image.
3.2 Applying SDS with Super-Resolution Diffusion Models
The main drawback of pixel-space diffusion models is that generating high-resolution images demands substantial computational resources. While latent diffusion models mitigate this cost by using a VAE [16] to encode images into a lower-dimensional latent space, pixel-space approaches typically rely on a cascaded generation process. In such pipelines, a diffusion model first produces a low-resolution image, and a subsequent super-resolution model upsamples it to the target resolution.
In [27], the authors apply SDS only with a low-resolution diffusion model. However, because this model is trained on downsampled images, it often fails to recover fine-grained details and high-frequency texture. To achieve high-fidelity outputs, we adopt the architectural structure of cascaded diffusion models. Specifically, we first run SDS with a low-resolution diffusion model, and then refine the output using SDS in combination with a super-resolution model. In the text-to-image setting, the second-stage model takes as input a low-resolution image along with a text prompt and timestep. This setup naturally suggests two strategies for applying SDS during refinement. The first strategy is to use the rendered image $g(\theta,c)$ as the diffusion-model input in SDS, while providing a downsampled version of the same image as a conditioning signal. Intuitively, this approach iteratively enriches the texture by adding details that were missing in previous gradient steps. The second strategy is to freeze the parameters $\theta_{0}=\theta$ obtained from the first stage, then run SDS with a fresh copy of $\theta$ as the optimization target and treat $\theta_{0}$ as a fixed conditioning input. In this case, the first-stage result serves as an anchor that guides the gradient descent dynamics.
In our experiments, the first strategy tends to be less stable and often drifts toward out-of-domain images, whereas the second strategy more reliably enhances resolution while preserving the details of the initial result. Figure 2 illustrates these observations.
@@ 3.3 Differentiable Physically Based Shading
Our texture synthesis approach applies SDS in a setting, where $g(\theta,z)$ renders an image of a fixed 3D model given a set of texture parameters $\theta$ and a random camera position and scene lighting specified by $z$. Rendering follows the differentiable rasterization pipeline of [25] implemented using Nvdiffrast [17].
Our textures follow the Disney physically based material model [4] with $\theta=(\theta_{d},\theta_{r},\theta_{m},\theta_{n})$ representing diffuse albedo, specular roughness, metalness, and a normal map, respectively. This parameterization ensures compatibility with standard real-time rendering engines. Lighting is provided via an environment map using a split-sum approximation of the outgoing radiance integral [14].
4 Score Distillation Sampling Analysis
In this section, we analyze Score Distillation Sampling (SDS) to motivate our proposed method. We argue that SDS applied to latent diffusion models is inherently prone to generating artifacts due to the ill-posed nature of the underlying optimization problem.